OpenStudy (anonymous):

Help please!!! Using Simpson's rule Determine the upper bound on the error in the following integral. Open to see it because i cant write equations in the top part of the question

OpenStudy (anonymous):

\[\int\limits_{1}^{3}(\sqrt[4]{X^2+1})\]

OpenStudy (anonymous):

This will be useful: https://en.wikipedia.org/wiki/Simpson%27s_rule#Error

OpenStudy (anonymous):

This means you'll have to find the upper bound of the fourth derivative of your given function. \[\begin{align*} f(x)&=(x^2+1)^{1/4}\\\\ f'(x)&=\frac{2x}{(x^2+1)^{3/4}}\\\\ f''(x)&=\frac{4x^2-3x+4}{2(x^2+1)^{7/4}}\\\\ f''''(x)&=\frac{-12x^3+15x^2-12x-6}{4(x^2+1)^{11/4}} \end{align*}\] (Check those derivatives, it's possible I made some mistake in the derivations)

OpenStudy (anonymous):

Definitely some mistakes...

OpenStudy (anonymous):

hmmm ya

OpenStudy (mathmate):

Correction: See also : http://openstudy.com/study#/updates/55a1e417e4b05670bbb52311 where \(f^{iv}(x)\) is given.

OpenStudy (anonymous):

eh I'm too lazy to solve it bai :3 btw the asker isn't here

OpenStudy (anonymous):

Hmm same problem, that's convenient :P

OpenStudy (anonymous):

xD ima help the guy with the salmon question